OpenStudy (anonymous):

Evaluate the integral x / (x^2 + 9) dx

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x }{ x^2 + 9 } dx\]

OpenStudy (helder_edwin):

use the substitution \[ \large u=x^2+9\qquad\qquad du=2x\,dx\quad\text{or}\quad \frac{du}{2}=x\,dx \]